I have a lot of trouble with word problems. Can someone help?
Beans sell for $9.00 a pound and $12.00 per pound. The two types need to be mixed to create a 100 pound mixture that sells for $11.25 per pound. How much of each bean should be used?
Thank you.

Answer by checkley75(3666)

100*11.25=12X+9(100-X)
1125=12X+900-9X
12X-9X=1125-900
3X=225
X=75 POUNDS OF $12.00 BEANS
100-75=25 POUNDS OF $9.00 BEANS.